Find the arithmetic mean an of the given terms.
1. an-1=0,an+1=5
2. an-1=2/3,an+1=4

Answer by stanbon(75887)

Find the arithmetic mean an of the given terms.
1. an-1=0,an+1=5
an+1 - an-1 = 2d
2d = 5-1 = 4
d = 2
----
an = 0+2
an = 2
-----------------------------------
2. an-1=2/3,an+1=4
an+1 - an-1 = 2d
2d = 4-(2/3) ] 10/3
d = 5/3
---
an = an-1 + d = (2/3) + 5/3 = 7/3
=======================================
Cheers,
Stan H.